### v1

# A Weak Penalty Neural ODE for Learning Chaotic Dynamics from Noisy Time Series

Abstract
Accurate forecasting of complex high-dimensional dynamical systems from observational data is essential for several applications across science and engineering. A key challenge, however, is that real-world measurements are often corrupted by noise, which severely degrades the performance of data-driven models. Particularly, in chaotic dynamical systems, where small errors amplify rapidly, it is challenging to identify a data-driven model from noisy data that achieves short-term accuracy while preserving long-term invariant properties. In this paper, we propose the use of the weak formulation as a complementary approach to the classical strong formulation of data-driven time-series forecasting models. Specifically, we focus on the neural ordinary differential equation (NODE) architecture. Unlike the standard strong formulation, which relies on the discretization of the NODE followed by optimization, the weak formulation constrains the model using a set of integrated residuals over temporal subdomains. While such a formulation yields an effective NODE model, we discover that the performance of a NODE can be further enhanced by employing this weak formulation as a penalty alongside the classical strong formulation-based learning. Through numerical demonstrations, we illustrate that our proposed training strategy, which we coined as the Weak-Penalty NODE (WP-NODE), achieves state-of-the-art forecasting accuracy and exceptional robustness across benchmark chaotic dynamical systems.
The accurate prediction of complex dynamical systems is of vital importance to science and engineering  [1, 2], with applications ranging from weather forecasting and climate modeling to fluid dynamics and systems biology. Traditional high-fidelity simulations are computationally expensive and often diverge from real-world data [3, 4]. Whereas data-driven methods [5, 6, 7, 8, 9], particularly those based on machine learning [10, 11, 12], have emerged as a powerful tool for learning these system dynamics directly from observational data and offer a computationally efficient alternative to first-principles simulation.
A variety of machine learning models have shown strong performance in forecasting complex spatiotemporal dynamics, including recurrent models  [13, 14] such as Long Short-Term Memory (LSTM) [15, 16, 14], reservoir computing (RCs) [14, 17, 18, 19], temporal convolutional networks (TCNs) [20, 21], attention-based Transformers [22, 23, 24], and more recently, random feature maps (RFMs) [25]. However, these models treat dynamics as a discrete-time autoregressive process, predicting each step from the last. This leads to error accumulation over long horizons and fails to capture invariant properties, such as probability density functions (PDFs), spectral characteristics, and attractors, limiting robustness to extended long-term forecasting. These issues are further exacerbated in the presence of observation noise.
Rather than learning discrete transitions, Neural Ordinary Differential Equations (Neural ODEs or NODEs) [26, 27] model dynamics by parameterizing the derivative of the state with a neural network. By leveraging modern ODE solvers, NODEs enable inference at arbitrary time points. This continuous formulation enables a flexible accuracy-speed trade-off and achieves substantially lower error than RNNs on time-series prediction [26]. Furthermore, NODEs offer architectural flexibility, with dynamics modeled by anything from simple feedforward networks to convolution [28], U-Net [28, 29], or transformer, enabling the capture of complex spatial and temporal dependencies beyond traditional recurrent models.
However, training NODEs remains challenging in practice. Standard Neural ODE training uses the adjoint sensitivity method [30, 31, 26], which, despite its memory efficiency, incurs high computational cost from repeated ODE forward and backward solves. Moreover, backward integration is often unstable, with exploding or vanishing gradients [32, 33, 34, 35], especially in stiff systems [36, 37]. Additionally, vanilla NODEs do not inherently guarantee stable long-horizon rollouts; achieving robustness typically requires additional architectural constraints or specialized training techniques [33, 32].
Most importantly, observational noise in real-world data severely limits the utility of this continuous-time framework [38, 39]. Pointwise losses force the model to fit noisy trajectories, making the learned dynamics overly sensitive to high-frequency fluctuations. This often degrades prediction accuracy and causes instability during inference, especially in chaotic systems where small errors can rapidly amplify into forecast failure [25, 40].
In response to the challenge of learning from noisy data, Bayesian and probabilistic models such as Gaussian Process-based approaches [41, 42, 43] and Bayesian neural surrogates [44, 45] offer principled uncertainty quantification, but often struggle with high computation cost, sensitivity to hyperparameters, restrictive modeling assumptions, and costly inference. Neural SDEs [46] extend probabilistic models by learning drift and diffusion terms, but focus on uncertainty modeling rather than recovering deterministic dynamics. The extended Kalman filter [47] estimates states recursively under noise but relies on known dynamics and linearization, which often fail in nonlinear settings. Notably, Galerkin-based methods employ weak formulations by projecting residuals onto test functions and minimizing integral errors, which naturally suppresses high-frequency noise. This weak-form spirit has inspired broader models [48, 49, 50], including weak-form or variational physics-informed neural networks (PINNs) [51, 52], weak-SINDy [40, 53], improving robustness by integrating residuals over subdomains. One recent work on weak-form NODEs [54] focuses on accelerating training and shows a promising speed-up, though its predictive accuracy does not consistently outperform baseline NODEs.
This reflects a key trade-off: integral nature provides robustness against noise but may potentially compromise predictive fidelity. Crucially, weak-form methods constrain dynamics only indirectly, bypassing the explicit forward simulation that serves as the ultimate test of a predictive model. In contrast, strong-form methods, though highly sensitive to noise, directly optimize pointwise residuals for high local fidelity. These complementary strengths and their respective shortcomings motivate a combined weak-strong training framework that reconciles the statistical stability of weak-form training with the local fidelity of strong-form supervision, especially under noisy conditions.
To mitigate this issue, we propose WP-NODE, a synergistic-form learning approach that combines weak and strong supervision within a NODE framework to robustly learn the dynamics of complex systems from noisy observational data. This method integrates the statistical robustness of weak-form training with the local accuracy of strong-form supervision, enabling the model to recover both the long-term statistical structure and the short-term dynamics of the system. Specifically, the training is guided by a unified loss function that combines shorter pointwise trajectory errors with weak-form residuals integrated against a set of test functions.
In the following sections, WP-NODE demonstrates its versatility and efficacy across several chaotic systems, including the 3-dimensional Lorenz-63 system, the 40-dimensional Lorenz-96 system, and the 64-dimensional Kuramoto–Sivashinsky equation. Our results show that this weak penalty formulation consistently improves both short-term predictive accuracy and long-term statistical consistency, offering a new perspective on robust data-driven modeling of complex dynamical systems.
I Results
Summary of Key Findings. The key results in this section can be summarized as follows. While weak-form supervision achieves relatively high short-term accuracy in local regions, its statistical consistency deteriorates, as revealed by significant invariant measure mismatches. In contrast, strong-form supervision captures the correct statistical distribution but suffers from poor short-term prediction accuracy due to training on noisy data. The proposed WP-NODE, however, balances both aspects, achieving accurate short-term prediction and faithful long-term statistical behavior.
Evaluation Metrics. To assess both short-term accuracy and long-term consistency, we use two complementary metrics: Valid Prediction Time (VPT) and Kullback–Leibler (KL) divergence [55]. The VPT is defined as the maximum duration (in units of Lyapunov time, $T_{\Lambda}=1/\Lambda$, where $\Lambda$ is the system’s maximal Lyapunov exponent), for which the normalized trajectory error remains below a certain error tolerance $\varepsilon$,
$\mathrm{VPT}=\frac{1}{T_{\Lambda}}\max\left\{n\Delta t\;\middle|\;E_{k}\leq\varepsilon\;\;\text{for all }k\leq n\right\}.$ (1)
Here, the normalized error (at time step $n$) is defined as $E_{n}=\sqrt{\frac{\sum_{j=1}^{D}\left(\hat{u}_{j}(n\Delta t)-u_{j}(n\Delta t)\right)^{2}}{\sum_{j=1}^{D}\sigma_{j}^{2}}}$. $D$ is the dimension of the system and $\sigma_{j}$ is the standard deviation of the $j$-th component of the true trajectory.
For long-term statistical consistency, the probability density function (PDF) of the surrogate model is estimated and compared to that of the reference system using the KL divergence. This comparison is performed dimension-wise via histogram-based approximations of the invariant measures. Visual comparisons of distributional statistics are also used to validate long-term behavior when KL divergence fails to reflect perceptual differences.
Lorenz-63. To illustrate the contrast between trajectory accuracy and statistical faithfulness, the 3D L63 [56] system is first considered. The governing equation is:
$\begin{aligned} \frac{dx}{dt}&=\sigma(y-x),\\ \frac{dy}{dt}&=x(\rho-z)-y,\\ \frac{dz}{dt}&=xy-\beta z\end{aligned}$ (2)
The well-known butterfly-like attractor can be reproduced under the standard parameters $\sigma=10.0$, $\rho=28.0$, and $\beta=\tfrac{8}{3}$. In our numerical experiment, we integrate this system of ODEs using the RK4 scheme with a time step of $\Delta t=0.01s$ for $100s$, yielding $N=10^{4}$ samples. Evaluation is based on a validation threshold of $\varepsilon=0.3$. Under these parameters, the largest Lyapunov exponent is $\Lambda\approx 0.91$.
First, $5\%$ data noise is applied to the observation data, where a zero-mean Gaussian is defined with standard deviation $\sigma=\sigma_{\mathrm{NR}}\cdot\mathrm{RMS}(x)$, based on the root-mean-square of the clean signal. The behavior of the trained WP-NODE model is illustrated in Fig. 1, demonstrating its great capabilities in both short-time prediction and attractor recovery. Specifically, Fig. 1a shows two representative trajectories corresponding to the best and worst predictions, selected from different initial conditions. The resulting VPT varies between 0.63 and 5.77 Lyapunov times. Despite the lower VPT in the second case, the prediction remains visually accurate for nearly 4 Lyapunov times before diverging. Figure 1b shows the corresponding Lorenz attractors, indicating that the WP-NODE robustly recovers the long-term statistical behavior of the system.
Figure 1: Forecasting of the Lorenz-63 system using WP-NODE under 5% training data noise. a. Time-series comparison between predicted and ground-truth trajectories for all three state variables. Rows show the best (VPT = 5.77) and worst (VPT = 0.63) cases (x-axis in Lyapunov time). Despite low VPT in the second case, predictions remain accurate for about 4 Lyapunov times. b. Reconstructed phase-space attractors for the predicted (orange) and true attractor (blue).
To complement the trajectory-based evaluation, Fig. 2 compares the predicted and ground-truth invariant measures using PDFs of the X component, for simplicity. These PDFs represent the system’s invariant measure across all methods and noise levels. Four methods, the proposed WP-NODE, and three baselines (strong NODE, weak NODE, and DeepSkip [25])are compared. All methods perform well under very low noise (1%). However, as noise increases to 5%, DeepSkip performance deteriorates, indicating high sensitivity to data perturbations. At higher noise levels (10% and 20%), both the strong and weak NODE also fail to recover the correct invariant measure, demonstrating a clear breakdown in statistical consistency under heavy noise. In contrast, WP-NODE remains robust, preserving a reasonable approximation of the true distribution even under 20% noise.
Figure 2: Invariant measure of the predicted Lorenz-63 system (for $100s$) for the X components beyond the training region, across different data noise. The proposed method maintains strong agreement with the ground truth across all noise levels, preserving long-term statistical properties. Other methods deteriorate significantly under moderate to high noise.
VPT (Lyapunov times) $\uparrow$ KL divergence $\downarrow$
System Method 0% 1% 5% 10% 20% 0% 1% 5% 10% 20%
[…]
Table 1: Cross-system evaluation (L63, L96, KS) under different observation noises. VPT is reported as mean $\pm$ std in Lyapunov times; KL divergence is reported as the mean value. The VPT evaluation is performed 30 times at random starting points beyond the training region. Higher is better for VPT, lower is better for KL divergence.
Figure 3: Performance comparison of WP-NODE and baselines models on the KS system, under 5% data noise. DeepSkip is not included since the original work was analyzed on a different configuration and would require task-specific fine-tuning. a. Short-time prediction comparison of the learned KS system under 5% observation noise. b. Joint probability density for the KS system under 5% observation noise. Both the WP-NODE and strong NODE closely reproduce the reference distribution, with WP-NODE showing slightly improved alignment in the core region. The weak NODE, however, exhibits notable distortion in the invariant measure.
Table 1 reports the VPT and KL divergence metric under varying noise levels (0%, 1%, 5%, 10%, and 20%), for all four methods. As shown on the left side of the table, for the L63 system, WP-NODE consistently outperforms both strong and weak NODE baselines under noisy conditions. This improvement can be attributed to two effects. When noise is low, the strong-form component provides stability and accuracy, enhancing short-term prediction. As noise increases, strong-form supervision becomes unreliable due to error amplification in pointwise residuals. In those regimes, the weak-form loss plays a dominant role in preserving local coherence. Notably, even at 20% noise, WP-NODE maintains high predictive accuracy, outperforming the weak form with a 14% gain in VPT. In contrast, DeepSkip, a leading baseline under clean data, deteriorates sharply with even small noise, highlighting the robustness of our proposed formulation.
Similarly, the right side of Table 1 reports the mean KL divergence metric as we defined above. While the KL metric tends to be small under low noise, due to the low-dimensional structure of the Lorenz attractor, WP-NODE continues to preserve lower values of KL, suggesting long-term statistical consistency, even at 10% and 20% noise levels. However, KL divergence becomes less informative at high values, as it may overlook localized distributional mismatches. To complement this, Supplementary Note 1 provides visual comparisons of the invariant measure (PDFs) for all three components across noise levels. These plots show that WP-NODE consistently preserves the true distribution, while baselines degrade significantly under high noise, highlighting the superior robustness and statistical fidelity of the proposed method.
To further evaluate robustness and generalizability, a comprehensive sensitivity analysis was conducted on the L63 system by varying key hyperparameters and training configurations. As shown in Supplementary Note 2, such a weak penalty formulation remains stable across a wide range of settings, highlighting its resilience to architectural and optimization choices.
[…]
Lorenz-96. To evaluate performance on high-dimensional chaotic dynamics, WP-NODE is tested on the 40D L96 model [57],
$\frac{dx_{i}}{dt}=(x_{i+1}-x_{i-2})x_{i-1}-x_{i}+F,\quad i=1,\dots,d,$ (3)
with cyclic boundary conditions $x_{i\pm kd}=x_{i}$ and forcing $F=10$. The system is integrated with step size $\Delta t=0.01$ for $1000$ seconds ($N=10^{5}$ steps). We measure the VPT with $\varepsilon=0.5$. This setup yields a leading Lyapunov exponent of $\Lambda\approx 1.68$ [25].
As shown in Table 1, all methods achieve comparable performance on this 40-dimensional L96 system, with only marginal differences in both VPT and KL divergence across noise levels. This similarity in performance may stem from the high-dimensional mixing of the system, which limits the discriminative power of standard metrics, or from inherent limitations of NODE-based methods in capturing fine-grained chaotic dynamics. Nonetheless, WP-NODE remains robust across all noise intensities, maintaining consistent predictive and statistical performance in this high-dimensional chaotic system. Additional comparisons and visualizations are provided in Supplementary Note 3.
Figure 4: Comparison of the long-term behavior of the KS system under 5% data noise, using the invariant measure (i.e., joint probability density). In the first row, WP-NODE shows more consistent results than the strong NODE in the second row, demonstrating its solver-agnostic property.
Kuramoto–Sivashinsky. The KS equation [58, 59] is a canonical benchmark for spatiotemporal chaos in nonlinear PDEs. It describes the evolution of a field $u(x,t)$ governed by
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}+\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{4}u}{\partial x^{4}}=0,$ (4)
where chaotic behavior emerges from the interaction between nonlinearity, instability, and dissipation.
The system is discretized using 64 Fourier modes on a periodic domain of length $22$, and integrated for $N=10^{5}$ steps with a time step of $\Delta t=0.25$, yielding a 25,000s25{,}000s trajectory. We measure the VPT with $\varepsilon=0.5$. In this regime, the largest Lyapunov exponent is approximately $\Lambda\approx 0.05$ [59].
Fig. 3a presents the short-time prediction skill of WP-NODE and related baselines on the KS system, where these models are trained with timeseries corrupted by 5% noise. We note that WP-NODE faithfully reproduces the dominant spatiotemporal structures, achieving a mean VPT of 3.31 (see Table 1). In contrast, strong and weak NODEs exhibit increasing phase drift and amplitude decay, with shorter predictive windows of 2.45 and 1.97, respectively. DeepSkip is excluded due to tuning difficulties and because the original work was trained on a different configuration [25].
Fig. 3b highlights WP-NODE’s ability to recover accurate long-term statistics via joint probability density plots. Under 5% observation noise, both WP-NODE and the strong NODE approximate the reference invariant distribution, but only WP-NODE captures both the core and tail behaviors with high fidelity. In contrast, the weak-form baseline fails to reproduce the full statistical structure, indicating degraded long-horizon dynamics.
Table 1 further reports robustness under different signal-to-noise ratios. WP-NODE maintains stable performance under perturbations, achieving a VPT of 2.87 at 10% noise, while the strong NODE baseline drops to 1.88. At higher noise levels, WP-NODE’s variant continues to outperform both baselines, highlighting its superior robustness. Interestingly, the weak NODE baseline shows erratic behavior: its VPT at 5% noise (1.97) is even lower than at 10% (2.42), indicating instability in dynamics reconstruction, consistent with prior findings [54]. In contrast, WP-NODE remains stable across noise levels. Overall, these results underscore a key strength of WP-NODE: it extends predictive horizons under noise while preserving the statistical properties of chaotic spatiotemporal systems, crucial for reliable long-term modeling.
Note that with 20% noise, WP-NODE utilizes 5 rollouts, which is found to be more effective for short-time prediction compared to lower-noise cases, where only a single rollout is used. This longer rollout strategy significantly improves VPT on the KS system, potentially because exposure to longer temporal contexts during training helps the model better distinguish signal from noise. In Supplementary Note 6, we provide the details of how many rollouts are used for each case in Table 1.
To further examine solver dependence, the trained model (using the dopri5 solver) was integrated under different numerical integration schemes, including bosh3, euler, midpoint, and rk4, for a $600s$ prediction (approximately 30 Lyapunov times) with 5% observation noise. As shown in Fig. 4, the strong NODE, despite being trained with 25-step rollouts, produces varying invariant measures across solvers, highlighting its sensitivity to the discretization scheme used during training. The weak NODE, lacking any rollout-based supervision, fails to capture long-term dynamics and becomes unstable across all solvers. Thus, we omit showing the joint probability density estimates from various ODE solvers. In contrast, the WP-NODE model produces density estimates that are not only qualitatively similar to the reference densities but also robust under different choices of ODE solvers. This empirical result suggests that the model has learned the underlying continuous-time dynamics in a solver-agnostic manner. This addresses a key limitation of purely weak or strong NODEs, whose predictive capability is typically either sensitive to time discretization or unstable during long-term integration.
II Conclusion and Discussion
This work introduced a weak penalty in training NODE, named WP-NODE, for learning chaotic dynamical systems from noisy data. Particularly, the proposed approach is to mitigate the lack of prediction skill of the standard NODE trained with a strong form formulation when the training data are corrupted by noise. Furthermore, the strong form NODE model’s long-term statistical prediction accuracy is sensitive to the choice of ODE integrator. While weak-form supervision alone achieves relatively high predictive horizons, it lacks statistical fidelity. We found that the proposed WP-NODE effectively reconciles both extended prediction time and robustly preserves long-term statistical structure. Across three chaotic systems, L63, L96, and KS, WP-NODE consistently delivers accurate short-term forecasts and preserves faithful long-term statistics, even under substantial observational noise.
[…]
Beyond efficiency, WP-NODE benefits from its combined weak–strong structure: the weak formulation naturally mitigates overfitting compared to pure pointwise losses, while the strong-form rollout enhances numerical stability by explicitly solving the underlying ODE. In many cases, the weak form alone trains successfully but fails at inference time due to exponentially exploding dynamics, making the weak penalty approach particularly effective for learning from noisy data.
However, the implemented weak formulation adopts predefined subdomain sizes and fixed test functions, which may influence performance in certain settings. While extensive ablation studies confirm robustness across a range of configurations, further improvements may be possible through adaptive subdomain partitioning, learnable test functions, or task-specific architectural tuning—particularly for more complex or higher-dimensional systems.
Overall, this study demonstrates that combining weak and strong supervision offers a simple yet effective route toward noise-robust, statistically consistent learning of chaotic dynamics, supporting both generalizability and extensibility to more complex systems.
III Method
Neural Ordinary Differential Equations. NODEs [26] generalize residual networks by replacing discrete-layer architectures with a continuous-time dynamics model. Given an initial condition $u(t_{0})=u_{0}$, the evolution of the state is governed by:
$\frac{du(t)}{dt}=f(u(t),t;\theta),$ (5)
where $f$ is a neural network parameterized by $\theta$. The solution at a later time $t$ is determined by solving this initial value problem using standard adaptive ODE solvers (e.g., Runge–Kutta, Dormand–Prince) [60, 61], which adjust their step size to balance computational cost and precision.
Strong Formulation and Adjoints. The standard training paradigm for a NODE is the strong formulation, which enforces that the predicted trajectory $\hat{u}(t)$ closely matches the ground-truth state $u(t)$. This objective is typically achieved by minimizing a loss function, such as the mean squared error (MSE),
$\mathcal{L}_{\text{strong}}(\theta)=\frac{1}{T}\sum_{n=0}^{T-1}\left\|u(t_{n})-\hat{u}(t_{n};\theta)\right\|^{2},$ (6)
where $T$ denotes the number of rollouts in training (set to $T=25$).
To optimize the parameters $\theta$ with respect to this loss, gradients are computed using the adjoint sensitivity method [26]. This approach provides exact gradients with constant memory cost, since intermediate forward states $u(t)$ do not need to be stored. However, the gradient computation becomes prohibitively expensive for chaotic or stiff dynamics, especially for large $T$. Exploding gradients and numerical instabilities demand small solver step sizes and tight tolerances, leading to high computational overhead [32, 36, 33]. These limitations motivate alternative strategies, such as weak-form or multi-step formulations, to improve stability and efficiency.
Weak Formulation. The weak form recasts the NODE into an integral form by weighting it with smooth and compactly supported test functions. This approach avoids pointwise enforcement of dynamics, improves robustness to noisy data, and eliminates the need for numerical differentiation in the training.
Let $\phi$ be a smooth test function defined on the time interval $[a,b]$, and satisfying $\phi(a)=\phi(b)=0$. The weak formulation of Eq. (5) requires that the residual is orthogonal to all test functions in an integral sense,
$\int_{a}^{b}\Bigl(\dot{u}(t)-f(u(t),t;\theta)\Bigr)\phi(t)dt=0.$ (7)
For numerical and theoretical advantages, we apply integration by parts to the derivative term $\int_{a}^{b}\phi(t)\dot{u}(t)dt$
$\int_{a}^{b}\phi(t)\dot{u}(t)dt=\Bigl[\phi(t)u(t)\Bigr]_{a}^{b}-\int_{a}^{b}\dot{\phi}(t)u(t)dt.$
The boundary contribution $\bigl[\phi(t)u(t)\bigr]_{a}^{b}$ vanishes because the test function satisfies $\phi(a)=\phi(b)=0$, leaving the weak formulation of Eq. (7) as
$\int_{a}^{b}\Bigl(u(t)\dot{\phi}(t)+f(u(t),t;\theta)\phi(t)\Bigr)dt=0,$ (8)
which avoids estimating $\dot{u}(t)$ from noisy data $u(t)$.
Subdomain Decomposition. To enhance flexibility and enable localized modeling, we employ the weak formulation to $K$ overlapping subdomains that cover the time domain of the available training time series. Each subdomain is mapped from the physical time coordinate $t\in[a_{k},b_{k}]$ to a local reference coordinate $s\in[-1,1]$ via a standard affine transformation $t=\tau_{k}(s):=\frac{b_{k}-a_{k}}{2}s+\frac{a_{k}+b_{k}}{2}$. Each subdomain contains $M$ uniformly spaced grid points (spacing $\Delta t$).
Weak Form Test Functions. The choice of test functions is critical for the accuracy and efficiency of weak-form integration. Following [49], we adopt the even polynomial family on the reference domain [1,1][-1,1],
$\phi_{p}(s)=(1-s^{2})^{p},\qquad s\in[-1,1],\;p\in\mathbb{N}.$ (9)
This construction ensures vanishing at the boundaries $(\phi_{p}(\pm 1)=0)$, allows smoothness to be tuned by the integer $p$, and, being polynomial, permits semi-analytical evaluation of the weak-form integrals. Detailed derivations are provided in the Supplementary Note 4.
Reference Mapping. Using the affine map $t=\tau_{k}(s)$, we have $ds=\tfrac{2}{L}\,dt$. Hence, $\frac{d\phi}{dt}=\frac{d\phi}{ds}\frac{ds}{dt}=\dot{\phi}(s)\frac{2}{L}$. The terms in the weak form from Eq. (8) transform as
$\begin{aligned} \int u(t)\dot{\phi}(t)dt&=\int_{-1}^{1}u(\tau_{k}(s))\dot{\phi}(s)ds,\\ \int f(u(t),t;\theta)\phi(t)dt&=\int_{-1}^{1}f(u(\tau_{k}(s)),\tau_{k}(s);\theta)\phi(s)\tfrac{L}{2}ds.\end{aligned}$
where $L=(M-1)\Delta t$ is the subdomain length. Overall, this localized formulation enables accurate and efficient evaluation of the weak residual within each subdomain using compactly supported test functions on [1,1][-1,1].
Weak Form Loss and Vectorization. To evaluate the weak form in practice, we represent both the data $u(s)$ and the function evaluations $f\big(u(s),s;\theta\big)$ using piecewise linear interpolation between $M$ grid points on each subdomain. This allows the weak form integrals to be expressed as linear combinations of the nodal values $u_{i}$ and the neural network evaluations at those nodes $f_{i}:=f(u_{i},t_{i};\theta)$. Particularly,
$\begin{aligned} \int u(\tau_{k}(s))\dot{\phi}(s)ds&\approx\sum_{i=0}^{M-1}u_{i}w_{\text{lhs},i}=\mathbf{u}^{\top}\mathbf{w}_{\text{lhs}},\\[6.0pt] \int f(u(\tau_{k}(s)),\tau_{k}(s);\theta)\phi(s)\tfrac{L}{2}ds&\approx\sum_{i=0}^{M-1}f_{i}w_{\text{rhs},i}=\mathbf{f}^{\top}\mathbf{w}_{\text{rhs}},\end{aligned}$
where the precomputable weights are $w_{\text{lhs},i}=\int\ell_{i}(s)\dot{\phi}(s)ds$ and $w_{\text{rhs},i}=\frac{L}{2}\int\ell_{i}(s)\phi(s)ds$. Here, $\ell_{i}(s)$ denotes the linear Lagrange basis function centered at the grid point $x_{i}$.
Since $\phi$ is smooth and $\ell_{i}(s)$ is linear, the weights admit exact closed-form expressions. In vectorized form, the weak-form loss is defined as the squared residual:
$\mathcal{L}_{\text{weak}}(\theta)=\big|\mathbf{u}^{\top}\mathbf{w}_{\text{lhs}}+\mathbf{f}^{\top}\mathbf{w}_{\text{rhs}}\big|^{2},$ (10)
where $\mathbf{f}$ denotes the vector of neural network evaluations at the nodal points $\mathbf{u}$ in each subdomain, i.e., $\mathbf{f}=\bigl[f(u_{0},t_{0};\theta),f(u_{1},t_{1};\theta),\dots,f(u_{M-1},t_{M-1};\theta)\bigr]^{\top}$. This formulation naturally generalizes to higher-dimensional or vector-valued data via batched dot products across dimensions. Detailed derivations, including the piecewise linear interpolation, the explicit form of the basis functions $\ell_{i}(s)$, and the construction of the weights, are provided in the Supplementary Note 5.
WP-NODE Loss Function. To leverage the complementary strengths of strong-form and weak-form supervision, WP-NODE is trained using a weak penalty added to the standard strong-form loss. The total loss is formulated as a weighted combination of Eq. (6) and Eq. (10), balancing pointwise accuracy with integrated residual consistency.
$\mathcal{L}=\mathcal{L}_{\text{weak}}+\lambda,\mathcal{L}_{\text{strong}},$ (11)
where $\lambda$ is a regularization coefficient balancing the two components.
A key distinction in our implementation is that $\mathcal{L}_{\text{strong}}$ is computed using significantly shorter forward-simulation rollouts (typically with $1\leq T\leq 5$) compared to standalone strong-form approaches. This design reflects a complementary role between the two loss components: The weak-form loss shapes the model’s global behavior via integral residuals while mitigating noise, whereas the strong-form loss provides a direct, localized supervisory signal. Such exposure to the true temporal evolution reinforces correct local dynamics, enhancing predictive fidelity without the cost or instability of long-horizon rollouts.
Dataset and Training. Training samples were generated using a sliding-window approach, where each windowed segment defines a short trajectory of fixed length, enabling consistent training across strong, weak, and WP-NODE formulations. Data is scaled using a MinMax scaler fitted on the training data. All neural network models used shallow neural networks (1–2 layers) with GELU activations and were trained with the Adam optimizer. A learning rate scheduler and early stopping based on validation loss are applied to ensure efficient convergence. For inference and reference solution generation, the adaptive dopri5 solver is employed in double precision with a strict numerical tolerance. Full details on all hyperparameters (e.g., subdomain sizes, batch sizes, learning rates, etc.) and other configurations are provided in Supplementary Note 6.
Acknowledgments This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility, and the Institute for Computational and Data Sciences (ICDS) at Pennsylvania State University. R.M. acknowledges support from the U.S. Army Research Office (ARO) Young Investigator Program under the Multiscale Modeling of Complex Systems, PM - Rob Martin. The research of J.H. was partially supported by the NSF Grant DMS-2505605 and the Office of Naval Research (ONR) grant N000142212193.
Author contributions J.H. and R.M. supervised the study. All authors contributed to the initial concept and machine learning methodology. X.L. carried out model development, implementation, training, and analysis. All authors contributed to writing the manuscript.
Competing interests The authors declare no competing interests.
Supplementary Information
I Supplementary Note 1: Lorenz 63 invariant measure analysis
To complement the quantitative KL divergence results presented in Table 1 of the article, this note provides a detailed visual comparison of the predicted and ground-truth PDFs for the L63 system under different modeling approaches. These comparisons offer a more intuitive understanding of how well each method captures the system’s invariant measure, especially under various noise levels.
As shown in Supplementary Fig. 1, each row corresponds to a different modeling approach (WP-NODE, strong NODE, weak NODE, and DeepSkip), and the columns show the estimated PDFs for the $x$, $y$, and $z$ state variables under 1%, 5%, 10%, and 20% noise levels. Overall, as the noise level increases, the mismatch between the predicted and true distributions becomes more significant for all methods. Notably, under 10% and 20% noise, both the weak and strong NODE struggle to recover meaningful distributions, often collapsing toward narrow or overly concentrated profiles that deviate substantially from the true attractor. In contrast, the proposed method continues to produce PDFs that closely resemble the ground truth across all state dimensions, demonstrating strong robustness in preserving long-term statistical structure even in highly noisy conditions.
These results reinforce the KL divergence findings and further highlight the advantage of the proposed method in capturing the correct invariant measure under real-world noise.
Supplementary Figure 1: Invariant measure comparisons of the learned Lorenz–63 system across different methods under varying noise conditions.
II Supplementary Note 2: Ablation study
A comprehensive ablation analysis was performed to investigate the influence of key hyperparameters. The default configuration consists of a two-layer neural network with 200 neurons per layer, subdomain size $M=60$, training signal length of $100s$, batch size of 1024, and number of subdomains $K=5000$ (half of the total sample size $N$). The polynomial order of the test function is $p=16$, the number of rollout trajectories is set to 1, and the regularization coefficient $\lambda=0.5$ balances the weak- and strong-form losses.
In Supplementary Fig. 2, each hyperparameter is varied individually while keeping all others fixed. Model performance is evaluated using the VPT and KL divergence metrics. As illustrated in Supplementary Fig. 2a, increasing the number of hidden layers leads to a noticeable degradation in performance. This counterintuitive trend likely arises from the higher sensitivity of deeper networks to input perturbations, making them more difficult to train under noisy conditions. Additional tests (not shown) conducted at higher noise levels further confirm that shallower architectures (e.g., single-layer networks) consistently outperform deeper ones. In Supplementary Fig. 2b, the effect of the integration subdomain size $M$ does not exhibit a clear optimum. Larger subdomains generally enhance numerical stability but reduce local adaptivity, whereas smaller ones capture finer local structures at the expense of stability.
In Supplementary Fig. 2c, the weak NODE and WP-NODE maintain high VPT scores even with substantially shorter training signals. For instance, accurate short-time prediction performance is achieved with training signals as short as $12.5s$, indicating that minimal temporal information is sufficient to capture the system’s essential dynamics under these formulations. In Supplementary Fig. 2d, batch size plays a critical role in training stability. Smaller batches often lead to unstable learning and divergent dynamics during inference, occasionally causing integration failures due to exploding values. Larger batch sizes, in contrast, stabilize optimization and improve generalization.
In Supplementary Fig. 2e, the model performance varies with the number of subdomains $K$. Intermediate values, typically around half of the total samples (e.g., $K=5000$ when $K_{\text{total}}=10000$), yield the best balance between short-term prediction accuracy and long-term statistical stability. This configuration is therefore adopted as the default setting in the experiments for the other two chaotic systems, L96 and KS. In Supplementary Fig. 2f, the polynomial order $p$ of the test function strongly influences weak-form integration accuracy. Values around $p=16$ yield optimal performance, likely due to the balanced smoothness and localization of the basis functions.
Notably, Supplementary Fig. 2g shows that increasing the number of rollouts in WP-NODE enhances pointwise residual minimization and temporal consistency. However, higher rollout counts also increase computational cost, highlighting a trade-off between accuracy and efficiency. Lastly, in Supplementary Fig. 2h, the regularization coefficient $\lambda$ critically affects performance. Moderate values (e.g., $\lambda=0.5$) balance weak-form integration and pointwise errors, achieving the best short-term performance.
Supplementary Figure 2: Comprehensive ablation study of the learned L63 system. a. Number of hidden layers, each layer has 200 neurons. b. The integration domain size $M$. c. The training signal length in seconds. d. The batch size during training. e. The number of subdomains $K$. f. The polynomial order $p$ of the test function. g. The number of rollouts used in WP-NODE for pointwise residual minimization. h. The regularization coefficient $\lambda$ balances the loss from weak and strong formulations.
[…]
Supplementary Figure 3: Per-epoch wall-clock time (in seconds) for training different models on the L63 system. WP-NODE is shown with varying rollouts, from 1 to 5. The strong NODE uses 25 rollouts. All models use a 2-layer network architecture with 200 hidden neurons. Lower times indicate faster training.
III Supplementary Note 3: Lorenz 96 analysis
Since all methods exhibit broadly similar performance on the L96 system, with WP-NODE showing only a modest improvement in both predictive accuracy and statistical consistency, only its results are presented here for clarity.
Supplementary Fig. 4 illustrates the extrapolation performance of WP-NODE on the 40-dimensional L96 system under 5% observational noise. The three rows correspond to different initial conditions, representing the best, average, and worst cases, with VPT scores of 4.29, 3.31, and 1.97, respectively. Across all cases, WP-NODE accurately reproduces the dominant spatiotemporal structures and maintains phase coherence over multiple Lyapunov times.
Supplementary Fig. 5 compares the invariant measures of the learned and reference systems across all 40 dimensions. The predicted distributions closely align with the ground truth, confirming that WP-NODE not only yields stable short-term forecasts but also preserves the long-term statistical equilibrium of the system. The slight deviations in a few components likely stem from high-dimensional mixing and the limited discriminative sensitivity of one-dimensional marginal densities. Overall, WP-NODE ensures stable extrapolation and accurate recovery of statistical properties, even in high-dimensional chaotic dynamics.
Supplementary Figure 4: Forecasting of the Lorenz–96 system using WP-NODE under 5% training data noise. The three rows correspond to three different initial conditions for extrapolation, representing the best, average, and worst cases. From top to bottom, the VPT scores are 4.29, 3.31, and 1.97.
Supplementary Figure 5: Invariant measures of the learned Lorenz–96 system using WP-NODE under 5% training data noise. The 40 dimensions are presented respectively.
IV Supplementary Note 4: Polynomial Test Functions
This note provides the explicit derivative expansions, antiderivatives, and weight constructions for the polynomial test functions $\phi_{p}(s)=(1-s^{2})^{p}$ used in the weak-form integration. These functions vanish at the subdomain boundaries, $\phi_{p}(\pm 1)=0$, which makes them particularly well-suited for localized formulations.
The $d^{\text{th}}$ derivative of $\phi_{p}$ admits the closed-form expansion
$\phi_{p}^{(d)}(s)=\sum_{k=0}^{p}(-1)^{k}\binom{p}{k}\frac{(2p-2k)!}{(2p-2k-d)!}s^{2(p-k)-d}.$
Since $\phi_{p}^{(d)}(s)$ and $s\phi_{p}^{(d)}(s)$ are polynomials, their primitives also have closed-form expressions:
$\begin{aligned} \Phi_{p,d}(s)&=\int\phi_{p}^{(d)}(s)ds,\\ \Psi_{p,d}(s)&=\int s\phi_{p}^{(d)}(s)ds.\end{aligned}$
The finite weights used in the weak-form summation are constructed from differences of these antiderivatives evaluated at subdomain boundaries.
V Supplementary Note 5: Piecewise Linear Interpolation and Decoupling of Integral
This note provides the detailed derivation of the weight vectors $w_{\text{lhs},i}$ and $w_{\text{rhs},i}$ used in the discrete weak form. We begin by defining the interpolation basis, then show how the weak-form integrals decompose into weighted sums of nodal values.
To evaluate the weak form in practice, we work within the reference domain $s\in[-1,1]$ of each subdomain and discretize it with $M$ grid points $s_{i}$. Let $s_{0}<s_{1}<\dots<s_{M-1}$ be partition points on the interval $s_{0},s_{M-1}]$, and suppose we are given values $x_{i}=x(s_{i})$ and $f_{i}=f(x_{i},s_{i};\theta)$ for $i=0,\dots,M-1$.
1. Linear Basis Functions. Define the piecewise linear Lagrange basis functions $\ell_{i}(s)$,
$\ell_{i}(s)=\begin{cases}\displaystyle\frac{s-s_{i-1}}{s_{i}-s_{i-1}},&s\in[s_{i-1},s_{i}]\\ \displaystyle\frac{s_{i+1}-s}{s_{i+1}-s_{i}},&s\in[s_{i},s_{i+1}]\\ 0,&\text{otherwise}\end{cases}\quad\text{for }1\leq i\leq M-2.$
For the endpoints:
$\ell_{0}(s)=\begin{cases}\displaystyle\frac{s_{1}-s}{s_{1}-s_{0}},&s\in[s_{0},s_{1}]\\ 0,&\text{otherwise}\end{cases}\quad\quad\ell_{M-1}(s)=\begin{cases}\displaystyle\frac{s-s_{M-2}}{s_{M-1}-s_{M-2}},&s\in[s_{M-2},s_{M-1}]\\ 0,&\text{otherwise}\end{cases}$
2. Interpolant Construction. We define the linear interpolants of the functions $x(s)$ and $f(u(s),s;\theta)$ as:
$\tilde{x}(s)=\sum_{i=0}^{M-1}x_{i}\ell_{i}(s),\quad\quad\tilde{f}(s)=\sum_{i=0}^{M-1}f_{i}\ell_{i}(s).$ (S1)
These functions are continuous and piecewise linear. Note that if $x_{i},f_{i}\in\mathbb{R}^{D}$ are vectors and $\ell_{i}(s)\in\mathbb{R}$ is a scalar, the terms $x_{i}\ell_{i}(s)$ and $f_{i}\ell_{i}(s)$ represent the standard scalar-vector products.
3. Projection Against a Test Function. We approximate the following integrals using Eq. (S1) as follows,
$\frac{L}{2}\int_{s_{0}}^{s_{M-1}}f(x(s),s;\theta)\phi(s)ds\approx\frac{L}{2}\int_{s_{0}}^{s_{M-1}}\tilde{f}(s)\phi(s)ds=\sum_{i=0}^{M-1}f_{i}\left(\frac{L}{2}\int_{s_{0}}^{s_{M-1}}\ell_{i}(s)\phi(s)ds\right)=\sum_{i=0}^{M-1}f_{i}w_{\text{rhs},i},$
where the weights are defined as:
$w_{\text{rhs},i}=\frac{L}{2}\int_{s_{0}}^{s_{M-1}}\ell_{i}(s)\phi(s)ds.$
Similarly,
$\int_{s_{0}}^{s_{M-1}}x(s)\phi^{\prime}(s)ds\approx\int_{s_{0}}^{s_{M-1}}\tilde{x}(s)\phi^{\prime}(s)ds=\sum_{i=0}^{M-1}x_{i}\left(\int_{s_{0}}^{s_{M-1}}\ell_{i}(s)\phi^{\prime}(s)ds\right)=\sum_{i=0}^{M-1}x_{i}w_{\text{lhs},i},$
where
$w_{\text{lhs},i}:=\int_{s_{0}}^{s_{M-1}}\ell_{i}(s)\phi^{\prime}(s)ds.$
These weights can be precomputed for a fixed $\phi(s)$, independent of $x$.
4. Explicit Expression for Weights. For interior indices $1\leq i\leq M-2$, we have:
$\begin{aligned} w_{\text{rhs},i}&=\frac{L}{2}\int_{s_{i-1}}^{s_{i}}\frac{s-s_{i-1}}{s_{i}-s_{i-1}}\phi(s)ds+\frac{L}{2}\int_{s_{i}}^{s_{i+1}}\frac{s_{i+1}-s}{s_{i+1}-s_{i}}\phi(s)ds\\ &=\frac{L}{2(s_{i}-s_{i-1})}\left[\Psi(s_{i})-\Psi(s_{i-1})-s_{i-1}(\Phi(s_{i})-\Phi(s_{i-1}))\right]\\ &\quad+\frac{L}{2(s_{i+1}-s_{i})}\left[s_{i+1}(\Phi(s_{i+1})-\Phi(s_{i}))-\Psi(s_{i+1})+\Psi(s_{i})\right],\end{aligned}$
where $\Phi_{p,d}(s)=\!\int\!\phi_{p}^{(d)}(s)ds$ and $\Psi_{p,d}(s)=\!\int\!s\phi_{p}^{(d)}(s)ds$.
For the endpoints:
$\begin{aligned} w_{\text{rhs},0}&=\frac{L}{2}\int_{s_{0}}^{s_{1}}\frac{s_{1}-s}{s_{1}-s_{0}}\phi(s)ds=\frac{L}{2(s_{1}-s_{0})}\left[s_{1}(\Phi(s_{1})-\Phi(s_{0}))-\Psi(s_{1})+\Psi(s_{0})\right],\end{aligned}$
and,
$\begin{aligned} w_{\text{rhs},M-1}&=\frac{L}{2}\int_{s_{M-2}}^{s_{M-1}}\frac{s-s_{M-2}}{s_{M-1}-s_{M-2}}\phi(s)ds\\ &=\frac{L}{2(s_{M-1}-s_{M-2})}\left[\Psi(s_{M-1})-\Psi(s_{M-2})-s_{M-2}(\Phi(s_{M-1})-\Phi(s_{M-2}))\right].\end{aligned}$
The other weights can also be computed as follows:
$\begin{aligned} w_{\text{lhs},i}&=\int_{s_{i-1}}^{s_{i}}\frac{s-s_{i-1}}{s_{i}-s_{i-1}}\phi^{\prime}(s)ds+\int_{s_{i}}^{s_{i+1}}\frac{s_{i+1}-s}{s_{i+1}-s_{i}}\phi^{\prime}(s)ds\\[6.0pt] &=\Bigl[\tfrac{s-s_{i-1}}{s_{i}-s_{i-1}}\phi(s)\Bigr]_{s_{i-1}}^{s_{i}}-\frac{1}{s_{i}-s_{i-1}}\int_{s_{i-1}}^{s_{i}}\phi(s)ds+\Bigl[\tfrac{s_{i+1}-s}{s_{i+1}-s_{i}}\phi(s)\Bigr]_{s_{i}}^{s_{i+1}}+\frac{1}{s_{i+1}-s_{i}}\int_{s_{i}}^{s_{i+1}}\phi(s)ds\\[8.0pt] &=-\frac{s_{i-1}}{s_{i}-s_{i-1}}\bigl[\Phi(s_{i})-\Phi(s_{i-1})\bigr]+\frac{1}{s_{i}-s_{i-1}}\bigl[\Psi(s_{i})-\Psi(s_{i-1})\bigr]\\[4.0pt] &\quad+\frac{s_{i+1}}{s_{i+1}-s_{i}}\bigl[\Phi(s_{i+1})-\Phi(s_{i})\bigr]-\frac{1}{s_{i+1}-s_{i}}\bigl[\Psi(s_{i+1})-\Psi(s_{i})\bigr].\end{aligned}$
For the endpoints, we similarly have:
$\begin{aligned} w_{\text{lhs},0}^{\prime}&=\frac{s_{1}}{s_{1}-s_{0}}\bigl[\Phi(s_{1})-\Phi(s_{0})\bigr]\;-\;\frac{1}{s_{1}-s_{0}}\bigl[\Psi(s_{1})-\Psi(s_{0})\bigr],\\[6.0pt] w_{\text{lhs},M-1}^{\prime}&=-\frac{s_{M-2}}{s_{M-1}-s_{M-2}}\bigl[\Phi(s_{M-1})-\Phi(s_{M-2})\bigr]\;+\;\frac{1}{s_{M-1}-s_{M-2}}\bigl[\Psi(s_{M-1})-\Psi(s_{M-2})\bigr].\end{aligned}$
VI Supplementary Note 6: Implementation and training details
Dataset Configuration After noise injection, the full time-series data is segmented into training samples using a sliding-window approach. Each sample consists of a trajectory segment of fixed length $T$, and in the strong formulation setting, the initial condition $y_{0}\in\mathbb{R}^{D}$ is also included. Given a trajectory of length $N$, this yields NTN-T training examples of the form $\{y_{0}^{(i)},\mathbf{y}^{(i)}\}_{i=1}^{N-T}$, where $y_{0}^{(i)}=x_{i}$ and $\mathbf{y}^{(i)}=\{x_{i},x_{i+1},\ldots,x_{i+T-1}\}$. Trajectories are extracted with uniform stride to maintain temporal coherence, enabling efficient mini-batch training across strong, weak, and WP-NODE formulations.
Hyperparameters and Training Setup To support the different training objectives in the framework, the above dataset construction strategies are adopted depending on the loss formulation.
The key hyperparameter settings employed in the experiments are summarized below. Strong-form training uses fixed sequence lengths of $25$ across all systems, while the WP-NODE uses shorter rollouts between $1$ and $5$ steps to reduce gradient instability. Details on the number of rollouts used for each case presented in the main article can be referred to Supplementary Table 1. For training components involving a weak-form loss (i.e., the weak NODE and WP-NODE modes), data is organized into integration subdomains. The subdomain size $M$ is set to $60$ for L63, $80$ for L96, and $60$ for the KS system. The number of overlapping subdomains $K$ is chosen to be roughly half the total data length $N$, and the test function parameter $p$ is set to 16, as these values were found empirically to yield optimal performance. The model is trained on the full time-series, since the weak-form loss targets integrated residuals over subdomains rather than pointwise supervision, making conventional overfitting to trajectory data less directly applicable.
Network Architecture and Optimization. All neural networks use the GELU activation function due to its consistent performance across tasks. Shallow architectures with 1–2 hidden layers are adopted to balance capacity and training stability. Training is performed using the Adam optimizer with minibatches. A batch size of $1024$ ($2048$ for KS) is used across different ensemble runs. A carefully tuned learning rate scheduler and early stopping based on validation loss are used to ensure efficient training. A relatively large initial learning rate of $0.02$ accelerates early convergence and is automatically decayed to refine performance.
For weak NODE and WP-NODE losses, models are trained for up to 20,000 epochs. In contrast, strong-form models are trained for fewer epochs (150–300, based on system complexity), due to higher computational cost and faster convergence. During inference, double-precision arithmetic is employed alongside the dopri5 solver to compute all reference trajectories with double precision, ensuring sufficient accuracy for reliable evaluation.
Number of rollouts $\downarrow$
System 0% 1% 5% 10% 20%
L63 1 2 2 1 2
L96 1 1 2 2 1
KS 1 1 1 1 5
Supplementary Table 1: Number of rollouts used during WP-NODE training for each case in Table 1.